40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 5 23:28:52 2019, rana, kruthi, Summary, SUS, ETMX unstuck by shaking the stack 
    Reply  Tue Jul 9 08:33:31 2019, gautam, Summary, SUS, ETMX PIT bias voltage changed by ~1V 
Message ID: 14730     Entry time: Fri Jul 5 23:28:52 2019     Reply to this: 14736
Author: rana, kruthi 
Type: Summary 
Category: SUS 
Subject: ETMX unstuck by shaking the stack 

We unstuck ETMX by shaking the stack. Most effective was to apply large periodic human sized force to the north STACIS mounts.

At first, we noticed that the face OSEMs showed nearly zero variation.

We tried unsticking it through the usual ways of putting large excitations through AWG into the pit/yaw/side DOFs. This produced only ~0.2 microns of motion as seen by the OSEMs.

After the stack shake, we used the IFO ALIGN sliders to get the oplev beam back on the QPD.

The ETMX sensor trends observed before and after the earthquake are attached.

** plots deleted; SOMEONE, tried to take raster images and turn them into PDF as if this would somehow satisfy our vetor graphics requirement. Boo. lpots must be actual vector graphics PDF

ELOG V3.1.3-